The perimeter of a rectangular plot of land is 312 kilometers. If the length is 6 kilometers more than 5 times the width, what are the length and width?
Answer by fion6002(19)

Perimeter of a rectangle = 2 width + 2 length
let x be the width
length = 5x+6
2x+(5x+6)2 = 312
2x+10x+12=312
12x=300
x=25
length = 5(25)+6 = 131
Therefore the width is 25km, the length is 131 km